At 27 ºC a gas occupies a volume of 600 cm3. What volume does that gas occupy at
227 ºC, assuming that pressure remains constant?
Give the name of the gas law associated with this scenario.
1
Expert's answer
2020-12-03T13:52:39-0500

"\frac{V1}{T1} = \frac{V2}{T2}"


"\frac{600}{300} = \frac{V2}{500}"


V2 = 1000cm3

gas law - charles law